OpenStudy (anonymous):

there are 72 students in a speech contest. Yesterday, 1/4 of them gave their speeches. Today 2/3 of the remaining students gave their speeches. How many students still haven't given their speeches?

OpenStudy (anonymous):

27 student's?

OpenStudy (anonymous):

72 * 1/4 = 18 students gave their speech 72 - 18 = 54 54 * 2/3 = 36 also gave their speech 54 - 36 = 18 students